OpenStudy (anonymous):

p-2√p-15=0 solve for p. simplify using radicals as needed, express complex numbers in terms of i.

OpenStudy (anonymous):

We use quadratic formula p= {2+or - sqrt[4-4*(1)(-15)] }/ 2 p ={2+or - sqrt[64] }/ 2 p = {2+- 8]/2 p = 10/2 or p = -6/2 p = 5 or -3

OpenStudy (anonymous):

dang, apparently answer was 25. Thanks though!

OpenStudy (anonymous):

that is because it should be \[\sqrt{p}=5\]